# Bridging the Gulf of Envisioning: Cognitive Design Challenges in LLM Interfaces

### v1

@@ 8. Conclusion
Our work applies a cognitive framework to characterize the dynamics of prompt-based interfaces such as ChatGPT, highlighting the complexities of interacting with LLMs. While LLMs are capable of interpreting a vast range of queries, their very flexibility can pose challenges for users attempting to convey precise intentions. We identify and characterize a new kind of interaction gulf called the “gulf of envisioning,” which captures the challenge users face in successfully formulating their intentions to elicit the desired response from an LLM. This gulf is further nuanced by the capability gap – what intentions can the LLM perform, the language gap –how to say what is needed to the LLM, and the intentionality gap – what to expect and how to evaluate the generated output, all of which describe varying facets of human-LLM misalignments. By arguing that for LLM interfaces, intentions are the interactions, we provide design recommendations to support the process of envisioning in generative tasks.

### v2

@@ 8. Conclusion
Our work applies a cognitive framework to characterize the dynamics of prompt-based interfaces such as ChatGPT, highlighting the complexities of interacting with LLMs. While LLMs are capable of interpreting a vast range of queries, their very flexibility can pose challenges for users attempting to convey precise intentions. We identify and characterize a new kind of interaction gulf called the “gulf of envisioning,” which captures the challenge users face in successfully formulating their intentions to elicit the desired response from an LLM. This gulf is further identified by the capability gap – what intentions can the LLM perform, the instruction gap –how to say what is needed to the LLM, and the intentionality gap – what to expect and how to evaluate the generated output, all of which describe varying facets of human-LLM misalignments. By arguing that for LLM interfaces, “intentions are actions,” we provide design recommendations to support the process of envisioning with generative AI models.
Acknowledgements.
We thank the reviewers for their feedback on the paper. Subramonyam and Agrawala are supported through the AI Research Institutes program by the National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award #2229873#2229873\#2229873# 2229873 - National AI Institute for Exceptional Education. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, the Institute of Education Sciences, or the U.S. Department of Education.